find the height of the racket ball when it left the racket??

A racket ball is struck in such a way that it leaves the racket with a speed of 4.87m/s in the horizontal direction. When the ball hits the court, it is a distance of1.95m from the racket. Find the height of the racket ball when it left the racket.
Answer #1
so first we do x direction: x=x0+ v0x t + 1/2 a t^2 in x direction a=0
so 1.95=0+4.87*t + 0.5 * 0 * t^2
so t=1.95/4.87=0.4
now y direction y=y0 + v0y t + 1/2 a t^2
0=y0+0*.4 + 0.5* (-9.8) * .4^2
so y0=0.5*9.8*0.4^2=0.784 m
answered by: ALISSA